A butcher has some hamburger that is 24% fat and some that is 21% fat. How many pounds of each should be mixed to make 87 lb of hamburger that is 22% fat?
How do I set this up?

Answer by richwmiller(17219)

a+b=87,
0.24*a+0.21*b=0.22*87
a=87-b
0.24*(87-b)+0.21*b=19.14
20.88-0.24b+0.21*b=19.14
-0.03*b=-1.74
b=58
a=87-b
a=29 lbs at 24% b=58 lbs at 21%
check
0.24*29+0.21*58=0.22*87
6.96+12.18=19.14
19.14=19.14
ok